OpenStudy (anonymous):

Simplify 5^-4/ 5^3 please help and explain how you did it. i dont understand

OpenStudy (anonymous):

\[\huge \frac{b^m}{b^n}=b^{m-n}\]

OpenStudy (anonymous):

\[\huge \frac{5^{-4}}{5^3}=5^{-4-3}\]would be step one

OpenStudy (anonymous):

so the answer is 5^-1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[-4-3\neq -1\]

OpenStudy (anonymous):

ugh i dont understand. What does that that thingy mean, for starters?

OpenStudy (anonymous):

5^-4/ 5^3 this is actually fairly simple. you figure out the exponents first so its would then be 0.0016/ 125 then divide so it would be 0.0000128